# Navi-plus: Managing Ambiguous GUI Navigation Tasks with Follow-up Questions

### v1

@@ Abstract
Graphical user interfaces (GUI) automation agents are emerging as powerful tools, enabling humans to accomplish increasingly complex tasks on smart devices. However, users often inadvertently omit key information when conveying tasks, which hinders agent performance in the current agent paradigm that does not support immediate user intervention. To address this issue, we introduce a Self-Correction GUI Navigation task that incorporates interactive information completion capabilities within GUI agents. We developed the Navi-plus dataset with GUI follow-up question-answer pairs, alongside a Dual-Stream Trajectory Evaluation method to benchmark this new capability. Our results show that agents equipped with the ability to ask GUI follow-up questions can fully recover their performance when faced with ambiguous user tasks.
Navi-plus: Managing Ambiguous GUI Navigation Tasks with Follow-up Questions
@@ 1 Introduction
Figure 1: Overview of our proposed Self-Correction GUI Navigation task. Agent proactively asks for missing information when the task is ambiguous.
@@
This issue has motivated us to review the formulation of the current GUI navigation task. In this paper, we propose a novel task called Self-Correction GUI Navigation, which endows GUI agents with a new ability to handle the ambiguity of human instruction. The core idea is to add an "ASK" action in the agents’ action space, enabling it to engage in intermediate natural language interactions with human users by proposing follow-up questions (Figure 1).
@@
We then include fair and comprehensive evaluation metrics to benchmark GUI agents’ capability to complete the GUI navigation task when the task description is ambiguous. A Dual-Stream Trajectory Evaluation method is proposed to separately compute metrics for the operational actions and for the additional ASK action, allowing direct model comparison.
@@
Our experiments show that the missing information in task description can significantly harm GUI agents task success rate, but with the completion of ASK action, agents’ can restore more than 99.4% of performance (Figure 2). Besides, we find that modern MLLM-based GUI agents can seamlessly learn the capability of proposing follow-up questions, and achieve satisfactory performance with timing accuracy above 0.932 and content accuracy exceeding 0.807.
2 Methods
2.1 Self-Correction GUI Navigation Task
@@
When users describe tasks, it is possible that some key information be omitted. For example, when ordering oil paint online with the help of GUI agents, someone specified colors and sizes but forgot to mention the preferred delivery method due to unfamiliarity with the task flow, leading to ambiguity in the task description for the GUI automation agent.
@@
To address the practical challenge of ambiguous task input, we propose a novel Self-Correction GUI Navigation task, aiming to benchmark and facilitate GUI agent feasibility when facing ambiguous task inputs. GUI agents’ ability to correctly continue the task and to interact with human users to complete the missing information are the two main indications we consider. (See 3.1)
@@
Specifically, we add an ASK action to the model’s action space, and during the interaction between the model and the user, we provide a SAY action for the user to fill in the missing information. The interaction between the agent and the user will be logged into the agent’s context, providing information to continue task navigation.
2.2 Dataset Construction Pipeline
@@
To intentionally generate ambiguous GUI navigation task descriptions and agent’s follow-up questions for information completion, we develop this data construction process. We select AndroidControl (Li et al., 2024a) and Mind2Web (Deng et al., 2024a) as our data sources for they are collected by well-trained human annotators to ensure quality.
We start by generating low-level instructions for each step in the trajectories with InternVL2.5-26B (Chen et al., 2024c) to identify each action’s operational intention. DeepSeek-V3 (Liu et al., 2024a) are then hired to decide whether a step is an informative step or a transactional step (for definitions, see Appendix C). Finally, DeepSeek-V3 is prompted to output an ambiguous task description that excludes selected informative steps, as "Task" and "Ambiguous Task" shown in Figure 1. A QA pair simulating the agent’s follow-up question and the user’s answer is also generated for each informative step, as the conversation between the agent and human shows in Figure 1.
More details of our data construction pipeline are described in Appendix D. For prompt templates see Appendix I. For dataset statistics see Appendix H.
@@ 2.3 Dual-Stream Trajectory Evaluation
Adding a new ASK action into agent action space makes the inference process and evaluation metrics inherently different from the original GUI navigation task. Since the evaluation is performed offline with pre-defined screenshots and trajectories, false-positive action predictions can occur if the ASK action appears before its annotated position. This over-strict criterion results in a diminish in observed agent performance.
So, we propose the Dual-Stream Trajectory Evaluation method with two key changes as depicted in Figure 3: (1) An ASK action is considered correct within a full task trajectory if it appears at or before the annotated step position. (2) If agents ASK in advance, execute a second inference by adding the ASK QA pairs into context to recover the operational action, like Figure 1.
See Algorithm 1 for pseudo code.
@@ 5             Insert ASK QA pair into $E_{m}$ ;
9             Replace step $a_{k}$ into $a_{k}^{\prime}$ in $A$.
@@ 3.1 Evaluation Metrics
Follow-up Question Metrics We propose to consider the timing and content relevance when evaluating the Self-Correction GUI Navigation task. The timing score focuses on the exact matches of ASK actions, and multiple scales including Precision, False Positive Rate (FPR), and F1 are calculated. The content relevance is measured by calculating the Cosine Similarity (CosSim) and METEOR score of the matched ASK actions.
For metrics formula see Appendix F.
3.2 Baseline Models and Implementation Details
We use the original data splits from AndroidControl (Li et al., 2024a) and Mind2Web (Deng et al., 2024a). We perform fine-tuning with the Qwen2.5-VL-3B (Qwen, 2025) and the SpiritSight-Agent-8B-base model adopting LoRA (Hu et al., 2021). For a more detailed explanation see Appendix E.
4 Results and Discussion
Operations ASK Timing ASK Content
Task Setting SSR SR SSR before / after Precision FPR F1 CosSim Meteor
SpiritSight-Agent-8B
w/ Original
Task
67.6 24.2 65.7 / 70.2 - -
59.6 16.7 60.8 / 60.2
w/ Incomplete Task 88.2% 69.0% 92.5% / 85.8% - -
62.5 18.7 51.3 / 70.3
+ ASK 92.5% 77.3% 78.1% / 100.1% 0.463 0.091 0.454 0.807 0.732
67.2 23.1 64.2 / 68.9 0.935 0.005 0.603 0.807 0.732
+ Dual Eval 99.4% 95.5% 97.8% / 98.1% +0.472 -0.086 +0.149
Qwen2.5VL-3B
w/ Original
Task
55.4 10.0 56.0 / 55.9 - -
51.3 8.7 51.6 / 52.3
w/ Incomplete Task 92.6% 87.0% 92.1% / 93.6% - -
52.6 9.5 47.9 / 56.2
+ ASK 94.9% 95.0% 85.5% / 100.5% 0.574 0.055 0.487 0.832 0.750
55.1 10.6 54.9 / 56.2 0.947 0.004 0.585 0.832 0.750
+ Dual Eval 99.5% 106.0% 98.0% / 100.2% +0.373 -0.051 +0.098
Table 1: Comparison of agents’ performance on original and incomplete task descriptions, using ASK actions, and using dual-stream evaluation on AndroidControl-Naviplus dataset. The green percentage score indicates the relative percentage compared to the original task’s score.
Ambiguous task descriptions can significantly degrade agent performance. Table 1 reports the performance of GUI agents on the AndroidControl-Naviplus dataset, comparing results for original and incomplete tasks. With the generated ambiguous task descriptions, the performance of baseline models drops significantly. SpiritSight-Agent’s SR, for example, decreased by 31% compared to the full-task scenario. Extended results with more missing information and results on Mind2Web-Naviplus, are reported in Appendix G.
The Dual-Stream Trajectory Evaluation is a more reasonable amendment to conventional navigation metrics. The "SSR before/after" column in Table 1 reports the SSR before and after the annotated ASK steps. The SSR scores after ASK steps recover when ASK action is applied, indicating the missing information has been recollected. However, SSR scores before ASK steps decrease compared to the no ASK steps case rather than recovery, demonstrating that the original evaluation method failed to take early ASK into consideration. In contrast, by applying our proposed dual-stream evaluation method, the SSR before and after the ASK steps align more closely.
Adding GUI follow-up questions recovers agent performances. Table 1 also compares the performance of GUI agents using incomplete tasks, with the addition of ASK actions and with Dual Stream Evaluation. The first three columns of operation metrics show that incorporating GUI follow-up QA annotations for training leads to a 26.5% increase for SpiritSight-Agent and an 8% increase for Qwen2.5VL in SR. The performance of all baseline models recovered to over 99.4% of their original Step SR scores and achieved an average SR score of 100.8% with the dual-stream evaluation. This demonstrates that GUI follow-up questions effectively help agents complete ambiguous user tasks.
The ability to propose GUI follow-up questions is perfectly acquired by MLLM-based GUI agents. The ASK action’s timing performance, as shown in Table 1, indicates that the agents can effectively ask follow-up questions when needed. Additionally, the cosine similarity score for the ASK content exceeds 0.807, while the Meteor score is above 0.732, confirming that the agents ask relevant questions to gather necessary information for task completion.
5 Related Work
5.1 Conversational AI Agents
We propose an interactive agent task to aid GUI automation, exploring the capability of conversational agents in GUI scenarios.
5.2 Conversational web navigation
6 Conclusion
Through this work, we introduced a novel Self-Correction GUI Navigation task, enlightening the ability of GUI automation agents to natively interact and complete missing information when faced with ambiguous user tasks. Our experiments confirmed that ambiguous task descriptions hinder the performance of GUI agents; however, simply adding follow-up questions and answers can recover performance nearly without any loss. Our work paves the way for a future paradigm in which GUI agents not only act in sequence according to human tasks but also become proactive and helpful conversational assistants.
Limitations
Practiced only on offline GUI navigation datasets, Future works should evaluate on online benchmarks. Our current Naviplus dataset only involves offline GUI navigation datasets as sources. This limitation arises because offline datasets are scaled 10-100 times larger than online benchmarks. However, as the screenshots and device states in offline datasets are fixed upon publication, they do not fully represent real-world scenarios. Future works should explore how the Self-Correction GUI Navigation task performs on online benchmarks, taking into account factors such as dataset scale and cross-dataset generalization.
Practiced only on mobile and web platforms, Future works should involve more platforms. Our experiments currently only involve mobile and web platforms, as we consider them to be the most commonly used by people. Nevertheless, future work should explore the Self-Correction GUI Navigation task on other platforms, such as desktop operating systems.
Practiced only on ambiguous tasks with one step’s information missed, Future works should explore more situations. We proposed the Self-Correction GUI Navigation task and evaluated it based on the base case where only one key piece of information is omitted by the user in a single step. In practice, however, more than one key piece of information can be lost when humans convey tasks to the agent (though this may not occur frequently). These scenarios should be explored in future work.
Future work should also discuss how the Self-Correction GUI Navigation task can be integrated with the planning capability. We identified that our proposed Self-Correction GUI Navigation task can be seamlessly integrated with the navigation planning and reflection framework proposed by our concurrent work, CogAgent-V2. Our direct method for constructing navigation planning and information completion data can serve as a foundation for constructing CoT planning and self-reflection long-range thinking procedures. Although we have addressed its relevance in Section 5.2, more discussion and experiments should be included in future work.
Appendix A Ethical Considerations
GUI automation agents have significant societal, security, and privacy implications. On one hand, they can free humans from repetitive operational tasks on digital devices and enhance work efficiency. On the other hand, if they fall into malicious hands, GUI agents could be misused to bypass anti-fraud systems or manipulate software to achieve harmful or unintended results. Additionally, GUI agents may make mistakes while performing tasks, leading to unacceptable results or unwanted side effects. There is also the risk of private information being leaked if proper data collection regulations are not in place. For these reasons, GUI automation agents must be fully regulated to ensure their broader use serves the social good.
Appendix B Extanded Related Work
B.1 GUI Navigation Agents and Datasets
Abundant data resources have been gathered to motivate the comprehensive development of VLM-powered GUI agents. Datasets like SeeClick(Cheng et al., 2024), GUICourse(Chen et al., 2024b), AMEX(Chai et al., 2024), SpiritSight(Huang et al., 2024), FerretUI-v1(You et al., 2024), FerretUI-v2(Li et al., 2024b), OS-ATLAS(Wu et al., 2024), ShowUI(Lin et al., 2024), AguVis(Xu et al., 2024) have collected large-scale screenshots annotated with element content and locations to enhance agents’ GUI element grounding capability. Other works such as RICOSCA(Deka et al., 2017), Widget Captioning(Li et al., 2020), Screen2Words(Wang et al., 2021), ScreenQA(Hsiao et al., 2022), WebSRC(Chen et al., 2021), GUICourse(Chen et al., 2024b) have constructed screenshot-conditioned question-answering data to enhance the agents’ domain knowledge in GUI environments. Furthermore, AMEX(Chai et al., 2024), GUICourse(Chen et al., 2024b), SpiritSight(Huang et al., 2024) have annotated the functionality of GUI elements to enable smoother generalization from element-level tasks to navigation tasks. Studies like Mind2Web(Deng et al., 2024a), OmniAct(Kapoor et al., 2024), GUICourse(Chen et al., 2024b), MoTIF(Burns et al., 2022), AITW(Rawles et al., 2024), GUI-Odyssey(Lu et al., 2024), AMEX(Chai et al., 2024), AndroidControl(Li et al., 2024a), OS-Genesis(Sun et al., 2024) have annotated real GUI navigation trajectories on various platforms for training and benchmarking practical GUI automation agents. AITZ(Zhang et al., 2024), AutoGLM(Liu et al., 2024b), CogAgent-V2(Hong et al., 2024), InfiGUIAgent(Liu et al., 2025) have augmented agents’ task planning capabilities by using CoAT approaches and self-reflections to empower scaling at inference-time.
We continue to advance the development of GUI automation agents by addressing the unavoidable problem of ambiguous user task presentations, thereby enhancing the agents’ capabilities and flexibility in real-world application scenarios.
Appendix C Task Formulation
Following WebCanvas(Pan et al., 2024)’s symbol that a real-world GUI task episode $E$ is $(S,A,T,O)$, where:
  • $S$ is the state space indicating the device status,
  • $A$ is the action space,
  • $T:S\times A\rightarrow S$ is the deterministic transition function,
  • $O$ denotes the state observation space.
We first formulate a GUI navigation task as:
$\text{Navi}(E,\text{Task})\rightarrow\left\{\begin{array}[]{l}\{i_{1},i_{2},\dots,i_{n}\},\\ \{a_{1},a_{2},\dots,a_{n}\},\\ \{s_{0},s_{1},\dots,s_{n}\}\end{array}\right.$
where Task is the user’s high-level instruction guiding the task, $a_{i}\in A$, $s_{i}\in S$, and $i_{i}\in I$ is the low-level instruction corresponding to the actions at each step.
In practical GUI navigation tasks, the execution process consists of both informative steps, which provide choices and guide decision-making in high-level instructions, and transactional steps, which represent inevitable actions necessary for task completion. Informative steps introduce decision branches (e.g., selecting options in a form), while transactional steps ensure the smooth flow of the task by performing essential operations (e.g., clicking confirmation buttons, and closing windows).
If some key information is omitted in the Task (becomes Task’), some informative steps may not be feasible, thus making the whole task less likely to be successful. To address this, we extend the action space $A$ by introducing an ASK action, which enables the agent to inquire about missing or ambiguous information. This modification results in the formulated Self-Correction GUI Navigation Task:
$\text{Navi}(E,\text{Task'})\rightarrow\left\{\begin{array}[]{l}\{i_{1},i_{2},\dots,i_{n}\},\\ \{a_{1},a_{2},\dots,a_{n}\},\\ \{s_{0},s_{1},\dots,s_{n}\}\end{array}\right.$
where the action space $A$ is adjusted to incorporate the ASK action.
Appendix D Data Construction Details
Low-level Instruction Completion We start by generating low-level instructions for each step in the trajectories. This intention consists of an operational intention and an element description (e.g. the ’OK’ button or the ’plus’ button of the second product). The current action, along with a screenshot and the bounding box of the interacted element, is provided to InternVL2.5-26B(Chen et al., 2024c) to generate the low-level instructions for that action.
Informative Step Decision We then hired DeepSeek-V3(Liu et al., 2024a) to decide whether a step is an informative step or a transactional step. DeepSeek-V3 achieves a satisfactory accuracy rate when making the judgment, with 90% of the data passing human verification.
Formation of Ambiguous Tasks We present the full task description as a reference to DeepSeek-V3 and provide it with the informative steps to be removed. The model is prompted to output an ambiguous task description that excludes the selected informative steps while maintaining all other information and the original style. A QA pair simulating the agent’s follow-up question and the user’s answer is also generated for each informative step. In practice, the judgment of informative steps and the formation of ambiguous tasks are completed in one API call to minimize costs.
Appendix E Experimental Setup In Detail
E.1 Baseline Models

### v2

@@ Abstract
Graphical user interfaces (GUI) automation agents are emerging as powerful tools, enabling humans to accomplish increasingly complex tasks on smart devices. However, users often inadvertently omit key information when conveying tasks, which hinders agent performance in the current agent paradigm that does not support immediate user intervention. To address this issue, we introduce a Self-Supplement GUI Navigation task that incorporates interactive information completion capabilities within GUI agents. We developed the Navi-plus dataset with GUI follow-up question-answer pairs, alongside a Dual-Stream Trajectory Evaluation method to benchmark this new capability. Our results show that agents equipped with the ability to ask GUI follow-up questions can interact with human users and recover their performance when faced with ambiguous user tasks.
Navi-plus: Managing Ambiguous GUI Navigation Tasks
with Follow-up Questions
@@ 1 Introduction
Figure 1: Overview of our proposed Self-Supplement GUI Navigation task. Agent proactively asks for missing information when the task is ambiguous.
@@
Figure 3: GUI tasks have developed from GUI environment understanding, through single-step instruction execution and multi-step navigation, and are evolving into more proactive and helpful assistants in the digital world.
@@
This issue has motivated us to review the formulation of the current GUI navigation task. In this paper, we propose a novel task called Self-Supplement GUI Navigation, which endows GUI agents with a new ability to handle the ambiguity of human instruction. The core idea is to add an "ASK" action in the agents’ action space, enabling it to engage in intermediate natural language interactions with human users by proposing follow-up questions (Figure 1).
@@
We then include fair and comprehensive evaluation metrics to benchmark GUI agents’ capability to complete the GUI navigation task when the task description is ambiguous. A Dual-Stream Trajectory Evaluation method is proposed to separately compute metrics for the operational actions (e.g., Step Success Rate) and for the additional ASK action, allowing direct model comparison.
@@
Our experiments show that the missing information in task description can significantly harm GUI agents task success rate, but with the completion of ASK action, agents can restore up to 99.4% of performance (Figure 2). Moreover, we find that modern MLLM-based GUI agents can seamlessly learn to propose follow-up questions and achieve satisfactory performance, with a timing accuracy of up to 0.947 and a content similarity of up to 0.832. Through extensive experiments, we further demonstrate that both model scale and dataset size affect performance on the proposed self-supplement GUI navigation task.
2 Related Work
2.1 GUI Navigation Agents and Datasets
GUI navigation agents (GUI agents in short) are data-driven, end-to-end and pure-vision-based agent models (Qin et al., 2025) that comprehend human instructions, perceive the virtual environment, and automate operations in the UI world. Simply providing the task with one sentence, the agent that functioned by a single model captures the screenshots and directly output the action to be conducted with one inference. Standing on the shoulder of vision-language foundation models, GUI agents further improve their capability in (1) GUI environment understanding, (2) operational elements grounding, and (3) navigation task planning. See Appendix E for detailed discussion.
In previous GUI navigation datasets, however, task instructions are deliberately designed to be unambiguous to ensure a unique action path. In this paper, we continue to advance the development of GUI navigation agents by addressing the unavoidable problem of ambiguous user task presentations, thereby enhancing the agents’ capabilities and flexibility in real-world application scenarios. We aim for a proactive GUI assistant experience. (Figure 3).
2.2 ASK Augmentation for Conversational Agents
The capability of actively putting up clarifying questions by conversational agents have been explored in the field of embodied robots (Nguyen and III, 2019; Ramrakhya et al., 2025), information retrieval (Chi et al., 2024), and text-to-SQL generation (Wu et al., 2024a). However, these previous works enhance model capability by integrating user assistance and ask human to provide explicit next-step guidance, which are not identical to GUI agent’s convention that only keep the agent responsible for decision making and execution.
Our work proposes an interactive agent task that refines user intentions with ASK to aid GUI navigation, exploring the proactive information completing capability of conversational agents in GUI scenarios.
2.3 Conversational Web Navigation
Figure 4: Illustration of Naviplus Dataset’s construction pipeline.
3 Self-Supplement GUI Navigation Task
3.1 Preliminaries
Symbol Description
$\mathcal{S}$ State space – all observable GUI states
$\mathcal{A}$ Action space – atomic GUI actions (e.g., click, type, swipe)
$T:\mathcal{S}\!\times\!\mathcal{A}\!\to\!\mathcal{S}$ Transition function – maps a state–action pair $(s,a)$ to the next state $s^{\prime}$
$\pi:\text{Task}\times\mathcal{S}\!\to\!\mathcal{A}$ Agent policy – issues an action given the task description and current state
$E$ GUI-navigation episode – ordered sequence of state–action pairs in one run
Table 1: Notation for GUI navigation.
Let Task be a natural-language task specification such as “compose and send an email to Alice.”
Given a specific Task, a GUI-navigation episode with $n$ steps is
$E(\text{Task})\;\triangleq\;\bigl{\langle}(s_{0},a_{0}),\,\dots,\,(s_{n-1},a_{n-1})\bigr{\rangle},\\ s_{0}\!\in\!\mathcal{S},\;a_{i}\!\in\!\mathcal{A}.$
At each step $i$ the agent receives $(\text{Task},\,s_{i})$ and issues
$a_{i}\;=\;\pi(\text{Task},\,s_{i}),$
after which the environment applies the transition function
$s_{i+1}\;=\;T(s_{i},a_{i}).$
The episode terminates when the goal specified by the Task is achieved or after $n$ steps, whichever occurs first.
In practical GUI navigation, the execution sequence interleaves informative steps and transactional steps, depending on whether the current interaction introduces a branching decision or performs an indispensable operation:
Informative steps presents alternative choices and guides high-level decision-making, introducing decision branches. eg. selecting an item from a drop-down list.
Transactional steps performs an indispensable operation required for task progress, ensuring a smooth workflow. eg. clicking an “OK” button or closing a pop-up window.
3.2 Task Formulation
@@
When users describe tasks, it is possible that some key information be omitted. For example, when ordering oil paint online with the help of GUI agents, someone specified colors and sizes but forgot to mention the preferred delivery method due to unfamiliarity with the task flow, leading to ambiguity in the task description for the GUI navigation agent.
@@
To address the practical challenge of ambiguous task input, we propose a novel Self-Supplement GUI Navigation task, aiming to benchmark and facilitate GUI agents’ feasibility when facing ambiguous task inputs. GUI agents’ ability to correctly continue the task and to interact with human users to complete the missing information are the two main indications we consider. (See 5.2)
@@
Specifically, we add an ASK action to the model’s action space for asking GUI Follow-up Questions, and during the interaction between the model and the user, we provide a SAY action for the user to fill in the missing information. The interaction between the agent and the user will be logged into the agent’s context, providing information to continue task navigation.
A Self-Supplement GUI Navigation Episode is thus formally defined as
$E^{+}\!\bigl{(}\textbf{Task}^{\prime}\bigr{)}\;\triangleq\;\bigl{\langle}(s_{0},a_{0}),\,\dots,\,(s_{n-1},a_{n-1})\bigr{\rangle},\qquad\\ s_{0}\in\mathcal{S},\;a_{i}\in\mathcal{A}^{+}.$
where original task description lacks some key information and becomes $\textbf{Task}^{\prime}$, Original action space is augmented by a ASK action and becomes $\mathcal{A}^{+}=\mathcal{A}\cup\{\text{ASK}\}$.
4 Navi-plus Dataset
4.1 Data Construction Pipeline
@@
An overview of our data construction pipeline is illustrated in Figure 4. To intentionally generate ambiguous GUI navigation task descriptions and agent’s follow-up questions for information completion, we develop this data construction process. We select AndroidControl (Li et al., 2024a) and Mind2Web (Deng et al., 2024a) as our data sources for they are collected by well-trained human annotators to ensure quality, and they cover the two most widely used device platforms: Mobile and Web.
The data construction process consists of three key steps: (1) Low-level Instruction Completion, (2) Informative Step Decision, (2) Formation of Ambiguous Tasks, and is described below in detail. To ensure the quality and reproducibility of our data, we select powerful open-source LLMs such as InternVL2.5-26B (Chen et al., 2024c) and DeepSeek-V3 (Liu et al., 2024a). The prompts and model outputs are verified by human annotators to ensure they meet or exceed the quality of GPT-4o. For prompt templates see Appendix F. For qualitative examples, see Figure X.
Low-level Instruction Completion We start by generating low-level instructions for each step in the trajectories. This intention consists of an operational intention and an element description (e.g. the ’OK’ button or the ’plus’ button of the second product). The current action, along with a screenshot and the bounding box of the interacted element, is provided to InternVL2.5-26B to generate the low-level instructions for that action, as shown on the left side of Figure 4. (A special case here is AndroidControl dataset originally provides human annotated low-level instructions.)
Informative Step Decision As shown in the middle of Figure 4, we then hired DeepSeek-V3 to decide whether a step is an informative step or a transactional step following the definition described in Section preliminary. Once a step is marked as transactional, it will be neglected, while only the informative steps will be included to generate QA. DeepSeek-V3 achieves a satisfactory accuracy rate when making the judgment, with 90% of the data passing human verification.
Formation of Ambiguous Tasks Finally, we present the full task description as a reference to DeepSeek-V3 and provide it with the informative steps to be removed. The model is prompted to output an ambiguous task description that excludes the selected informative steps while maintaining all other information and the original phrasing style. A QA pair simulating the agent’s follow-up question and the user’s answer is also generated for each informative step, as shown on the right side of Figure 4. In practice, the judgment of informative steps and the formation of ambiguous tasks are completed in one API call to minimize costs.
5 Evaluation Methods
@@ 5.1 Dual-Stream Trajectory Evaluation
Adding a new ASK action into agent action space makes the inference process and evaluation metrics inherently different from the original GUI navigation task. Since the evaluation is performed offline with pre-defined screenshots and trajectories, false-positive action predictions can occur if the ASK action appears before its annotated position. This over-strict criterion results in a decline in observed agent performance, which requires correction.
So, we propose the Dual-Stream Trajectory Evaluation method with two key changes as depicted in Figure 5: (1) An ASK action is considered correct within a full task trajectory if it appears at or before the annotated step position. (2) If agents ASK in advance, execute a second inference by adding the ASK QA pairs into context to recover the operational action, like Figure 1 shows.
During the second inference pass, as illustrated in Figure 5, ASK QA pairs are inserted into prompts from the predicted position to its annotated position. This design ensures that once an ASK action is invoked, it affects all subsequent steps in the trajectory, reflecting a real-world execution scenario. Moreover, the proposed method isolates the assessment of the ASK action from the evaluation of other operational steps, ensuring that the model’s ability to ask questions and execute actions can be measured independently.
See Algorithm 1 for pseudo code, with an additional computational cost of O(n) for the second evaluation pass.
@@ 5      Insert ASK QA pair into $E_{m}$ ;
9      Replace step $a_{k}$ into $a_{k}^{\prime}$ in AA.
@@ 5.2 Evaluation Metrics
Follow-up Question Metrics We propose to consider the timing and content relevance when evaluating the Self-Supplement GUI Navigation task. The timing score focuses on the exact matches of ASK actions, and multiple scales including Precision, False Positive Rate (FPR), and F1 are calculated. The content relevance is measured by calculating the Cosine Similarity (CosSim [^1] Cosine similarity of Sentence Transformer embedding.) and METEOR score of the matched ASK actions.
For metrics formula see Appendix D.
6 Experimental Setup
6.1 Baseline Models